Round to two decimal places as needed 9 K.
Answer #1

Po=0.5
o-1 Po 0.5
n = 808
P = X / n = 501/808 = 0.62


Test Statistic :-
Z = ( P - P_{0}) / ( \sqrt{(P_{0} * q_{0})/n})
Z = (0.62-0.5)/(V/( 0.5 0.5)/808 )
Z = 6.82